# FE-LWS: Refined Image-Text Representations via Decoder Stacking and Fused Encodings for Remote Sensing Image Captioning

### v2

@@ Abstract
Remote sensing image captioning (RSIC) aims to generate meaningful textual descriptions from remote sensing images using encoder-decoder frameworks. Although previous efforts have focused primarily on improving the design of decoders, significant challenges persist in capturing the complex spatial and semantic characteristics of remote sensing data. In addition, reliance on a single encoder or decoder often limits the representational capacity of the model. To address these limitations, we propose a novel framework that fuses features from two complementary CNN encoders and integrates a stacked GRU decoder with weighted averaging to better model multilevel semantics. Experimental evaluations demonstrate that the proposed approach outperforms existing baselines, delivering more accurate and contextually rich captions for remote sensing images.
Index Terms:
Remote Sensing Image Captioning (RSIC), Encoder-Decoder model, Convolutional Neural Network (CNN), Gated Recurrent Unit (GRU), Fusion-based Encoders (FE), Weighted-based Stacking (WS).
I Introduction
Remote Sensing Image Captioning (RSIC) is a task in the machine learning domain that focuses on extracting meaningful information from a remote sensing images. The encoder-decoder framework is the dominant approach in RSIC, where a CNN-based encoder (such as ResNet [1], ConvNext [2]) extracts image features and a decoder (e.g., LSTM [3, 4], GRU [5], or SVM [6]) generates captions in sequence-to-sequence manner.
Although decoder optimization has received considerable attention [3, 7], the encoder’s role in capturing spatial and semantic diversity remains underexplored. Most of the methods are based on a single CNN, which limits the representational richness. To address this, we propose a fusion-based encoder that integrates features from two complementary CNNs, enabling a more detailed and robust image understanding.
Similarly, most existing models employ a single decoder, which restricts the modeling of hierarchical semantics and long-range dependencies [4, 8, 9]. We adopt a stacked GRU-based decoder [10, 11] to sequentially refine features across layers, improve contextual representation, and mitigate vanishing gradients. To avoid losing intermediate information, we aggregated the outputs of all decoder layers using a weighted averaging strategy, allowing the model to capture both low and high-level semantics more effectively.
The major contributions of this letter are as follows.
  • A fusion-based encoding strategy is proposed to enhance image representation and understanding.
@@ I Introduction
  • A novel multi-layer stacked GRU decoder with weighted averaging is proposed to improve sequence modeling.
  • The proposed model is thoroughly evaluated against state-of-the-art methods, demonstrating its notable performance improvement.
The structure of the letter is as follows: Section II introduces our proposed approach, focusing on fusion-based encoding and stacking-based decoding techniques. Section III describes the experimental setup, covering datasets, model configurations, and evaluation metrics. Table IX presents the results of various aspects, including ablation studies of our model, and provides a comparative analysis with baseline models through both quantitative and qualitative evaluations, supported by visual examples. Finally, Section V summarizes the key findings of the letter and highlights potential avenues for future research.
II Proposed Method
@@
The overall architecture is illustrated in Figure 1. We adopt an encoder-decoder framework, where the encoder ($F$) consists of two distinct CNNs whose outputs are concatenated and processed through a linear transformation layer ($L_{1}$). For decoding, a tokenized input sequence passes through an embedding layer ($E$) and then into a stack of Gated Recurrent Units (GRU) ($G$) to model temporal dependencies. The outputs from $L_{1}$ and $G$ are concatenated and passed through two additional linear layers ($L_{2}$ and $L_{3}$), with $L_{3}$ serving as a softmax layer to generate word probabilities.
II-A Fusion of Different Encoders
Figure 2: Architecture of the Fusion Technique Used in Our Model
The proposed method integrates features extracted from two widely used CNN architectures. The fusion strategy is illustrated in Figure 2. Given the vast number of CNN models available, an exhaustive comparison of all individual models and their pairwise combinations is impractical. Therefore, this study focuses on four prominent architectures: ConvNext [2], ResNet [1], ResNext [12], and DenseNet [13]. These models have demonstrated strong feature extraction capabilities in RSIC [4, 14].
TABLE I: Performance Comparison of Individual and Fusion-Based CNN Encoders on the SYDNEY Dataset
@@ II-A Fusion of Different Encoders
TABLE II: Performance Comparison of Individual and Fusion-Based CNN Encoders on the UCM Dataset
@@
TABLE III: Performance Comparison of Individual and Fusion-Based CNN Encoders on the RSICD Dataset
@@
As shown in Tables III, III and III, the combination of features from two different CNNs improves performance compared to the use of a single CNN. This fusion does not add significant computational cost, as CNNs are used solely for feature extraction and remain frozen during training. Since the encoder parameters are not updated, the training remains efficient while benefiting from the diverse spatial and semantic information captured by multiple CNNs. Among the combinations, ConvNext and ResNet yield the best results due to their complementary properties. ConvNext effectively captures global spatial patterns, while ResNet retains fine-grained details through residual connections. Previous studies [4, 14] also highlight their superiority in RSIC tasks. Their fusion results in richer and more robust feature representations, making ConvNext and ResNet the chosen pair for all subsequent experiments in this study.
II-B Stacking of Multiple Decoders
Our model uses the Gated Recurrent Unit (GRU) [15] as the decoder. The GRU effectively manages the flow of information through its gating mechanisms, allowing it to maintain crucial temporal dependencies while minimizing computational complexity. Conventional encoder-decoder architectures typically rely on a single decoder to generate captions in a sequence-to-sequence manner [16, 7, 6, 4]. However, recent work has shown that stacking multiple decoders enhances contextual modeling and hierarchical learning by refining intermediate representations and improving long-term dependency capture [17, 18, 19]. In our approach, we employ a stacked GRU structure in the decoder to enhance the model’s ability to process sequential data and to improve its robustness in RSIC. However, simply stacking GRUs sequentially and relying solely on the final layer’s output does not always yield the optimal result because deeper networks can suffer from vanishing-gradient problems [20, 21], and important intermediate features might be lost if only the last GRU’s output is considered. Furthermore, the information flow between stacking layers may not be efficiently balanced, leading to suboptimal learning. To address these challenges, we explore four different stacking strategies.
TABLE IV: Stacked Decoder Results on SYDNEY Dataset
@@ II-B Stacking of Multiple Decoders
TABLE V: Stacked Decoder Results on UCM Dataset
@@
TABLE VI: Stacked Decoder Results on RSICD Dataset
@@
II-B1 Simple Stacking
The traditional stacking approach, called Simple Stacking (SS), directly utilizes the output of the final decoder layer as the final output of the system. The overall output of the SS system is given below.
$O=D_{n}$ (1)
II-B2 Concatenation-based Stacking
In Concatenation-based Stacking (CS), the outputs of all decoder layers are concatenated to form the final output, thus preserving a richer set of feature representations. Then, the overall output $O$ of the CS system is given below.
$O=D_{1}|D_{2}|...|D_{n}$ (2)
II-B3 Global Weighted-based Stacking
In Global Weighted-based Stacking (GWS), a weighted average of the output of all decoder layers is calculated using a single global weight, ensuring a balanced contribution from each layer. The output of the GWS system is given below.
$O=\sum_{i=1}^{n}\alpha_{i}{D_{i}}$ (3)
II-B4 Local Weighted-based Stacking
In Local Weighted-based Stacking (LWS), a weighted average of the decoder output is calculated using element-specific weights, allowing for finer control over the contribution of each feature to the final output. The output of the LWS system is given below.
$O=\sum_{i=1}^{n}W_{i}\odot D_{i}$ (4)
Here, $n$ denotes the number of decoders, and $D_{i}$ represents the output of the $i^{\text{th}}$ decoder, while $O$ is the final output of the stacked system. The symbol |||| indicates the concatenation operation. The parameter $\alpha_{i}$ represents a scalar weight, while $W_{i}$ is a vector weight of the same size as $D_{i}$ for each $D_{i}$. Furthermore, $0\leq\alpha_{i}\leq 1$ with constraint $\sum_{i=1}^{n}\alpha_{i}=1$, and $0\leq W_{i}^{j}\leq 1$ with $\sum_{i=1}^{n}W_{i}^{j}=1$, where $W_{i}^{j}$ denotes the value of the characteristic $j^{\text{th}}$ of $W_{i}$. Both $\alpha=\{\alpha_{i}\mid i=1,\dots,n\}$ and $W=\{W_{i}\mid i=1,\dots,n\}$ are trainable parameters.
The impact of the depth of the decoder is evaluated in Tables VI, VI and VI, where the number of stacked GRUs (DC) is treated as a hyperparameter ranging from two to five. Configurations beyond five are excluded due to increased computational cost and diminishing returns.
In the SYDNEY and UCM datasets( Tables VI and VI), deeper stacks under SS and CS show inconsistent trends, often degrading beyond two layers. SS, in particular, underperforms the single-decoder baseline (NS), probably due to vanishing gradients and limited training data [22]. RSICD maintains stable performance up to three layers, with declines beyond that point, indicating sensitivity to the depth of the decoder.
From the results, we observe that among the stacking strategies, SS and CS are less effective. SS eliminates intermediate outputs, while CS introduces redundancy by treating all outputs equally. Weighted strategies offer a better balance: GWS uses global weights but lacks temporal flexibility, whereas LWS applies adaptive element-wise weights, enabling finer control over temporal dependencies. LWS consistently achieves the best results, with three stacked decoders yielding optimal performance across all datasets.
III Experimental Setup
We use a conventional encoder-decoder architecture. Each GRU on the decoder stack ($G$) has an output size of $256$. Linear layers $L1$, $L2$, and $L3$ have output sizes of $256$, $512$, and one more than the vocabulary size, respectively. GELU is used as the activation function for $L1$ and $L2$, while $L3$ uses SoftMax. A dropout rate of $0.5$ is applied before $L1$, after the embedding layer ($E$), before the sequence loss layer ($SL$), and between consecutive GRUs to prevent overfitting. The embedding size is set to $256$.
@@ III Experimental Setup
The experiments were carried out using TensorFlow within a Docker environment with GPU support, on a host machine equipped with a NVIDIA RTX A6000 GPU with $50GB$ memory and $132GB$ RAM. During manuscript preparation. ChatGPT and Grammarly were used solely for rephrasing and language refinement, without generating any original technical content.
III-A Datasets Used
The performance of the model was evaluated using three widely recognized RSIC datasets: SYDNEY [16], UCM [16], and RSICD [22]. The details of these datasets are provided in the supplementary material. In this study, we used the corrected versions of these datasets [3], which address common issues such as spelling mistakes, grammatical errors, and inconsistent dialect variations of English. We used the same train-validation-test split as defined in these datasets.
III-B Performance metrics
The evaluation metrics used in our model include BLEU [23], METEOR [24], ROUGE [25], and CIDEr [26], which are widely adopted in RSIC tasks [16, 22, 3, 4]. More details are provided in the supplementary material.
IV Results
TABLE VII: Performance Comparison of Model Variants Through Ablation Studies on SYDNEY Dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.7531 0.6645 0.5982 0.5376 0.3750 0.6951 2.3954
R-NS 0.7363 0.6502 0.5785 0.5138 0.3621 0.6809 2.2526
FE-NS 0.7679 0.6814 0.6084 0.5458 0.3801 0.7263 2.4515
C-LWS 0.8141 0.7321 0.6610 0.6054 0.4369 0.7584 2.7472
R-LWS 0.8056 0.7211 0.6527 0.5924 0.4216 0.7364 2.5743
FE-LWS 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
TABLE VIII: Performance Comparison of Model Variants Through Ablation Studies on UCM dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.8369 0.7804 0.7312 0.6853 0.4689 0.8037 3.3990
R-NS 0.8297 0.7705 0.7178 0.6683 0.4506 0.7903 3.3329
FE-NS 0.8457 0.7919 0.7533 0.7166 0.4833 0.8163 3.4255
C-LWS 0.8655 0.8121 0.7644 0.7192 0.5086 0.8367 3.4958
R-LWS 0.8614 0.8034 0.7526 0.7049 0.4910 0.8194 3.4621
FE-LWS 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
TABLE IX: Performance Comparison of Model Variants Through Ablation Studies on RSICD dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.6327 0.4636 0.3602 0.2903 0.2487 0.4648 0.8272
R-NS 0.6322 0.4621 0.3557 0.2835 0.2489 0.4649 0.8206
FE-NS 0.6452 0.4767 0.3732 0.3087 0.2609 0.4827 0.8594
C-LWS 0.6671 0.5008 0.3996 0.3302 0.2843 0.5066 0.8954
R-LWS 0.6645 0.4981 0.3963 0.3257 0.2849 0.5010 0.8684
FE-LWS 0.6796 0.5124 0.4073 0.3365 0.2944 0.5248 0.9423
This section presents a comprehensive evaluation of the proposed model through ablation studies, quantitative comparisons with baseline methods, and visual analysis. The goal is to assess the individual and combined contributions of encoder fusion, decoder stacking, and caption selection strategy.
IV-A Ablation Studies
To evaluate the contribution of individual components, we compare four encoder-decoder configurations: single encoder with single decoder, fusion-based encoder with single decoder, single encoder with LWS decoder and the proposed fusion-based encoder with LWS decoder.
The results, shown in Tables IX, IX and IX, follow the naming convention where C denotes ConvNext,R denotes ResNet, andFE indicates their fusion. For decoders,NS refers to a single decoder andLWS to Local Weighted-based Stacking with three GRUs (see Section II-B).
The results indicate that both encoder fusion and decoder stacking improve performance, with the FE-LWS configuration achieving the highest scores across all datasets. In particular, LWS offers greater performance gains than encoder fusion alone, highlighting its effectiveness even when used with a single CNN.
IV-B Numerical Evaluation
TABLE X: Experimental Results of Baseline Methods and Different Searching Techniques on SYDNEY Dataset

### v3

# MsEdF: A Multi-stream Encoder-decoder Framework for Remote Sensing Image Captioning

@@ Abstract
Remote sensing images contain complex spatial patterns and semantic structures, which makes the captioning model difficult to accurately describe. Encoder-decoder architectures have become the widely used approach for RSIC by translating visual content into descriptive text. However, many existing methods rely on a single-stream architecture, which weakens the model to accurately describe the image. Such single-stream architectures typically struggle to extract diverse spatial features or capture complex semantic relationships, limiting their effectiveness in scenes with high intraclass similarity or contextual ambiguity. In this work, we propose a novel Multi-stream Encoder-decoder Framework (MsEdF) which improves the performance of RSIC by optimizing both the spatial representation and language generation of encoder-decoder architecture. The encoder fuses information from two complementary image encoders, thereby promoting feature diversity through the integration of multiscale and structurally distinct cues. To improve the capture of context-aware descriptions, we refine the input sequence’s semantic modeling on the decoder side using a stacked GRU architecture with an element-wise aggregation scheme. Experiments on three benchmark RSIC datasets show that MsEdF outperforms several baseline models.
1 Introduction
Remote sensing image captioning (RSIC) refers to the task of automatically generating descriptive sentences from high‑resolution remote sensing images. These images often contain complex spatial layouts, overlapping semantic objects, and ambiguous visual features that challenge traditional representation methods. A widely used approach is the encoder and decoder framework, where a convolutional neural network (CNN) extracts semantic features and a sequential model, typically an RNN or LSTM, generates textual descriptions [18, 16]. To improve performance, previous work has used fully convolutional encoders [20], summarization-driven training [21], and SVM based decoders [7]. More recent efforts include improving the decoder embedding layer with TextGCN embeddings [4] and dynamic graph refinement using CLIP-based semantic and spatial fusion [14].
Most existing approaches for RSIC still rely on conventional single-stream encoder and decoder architectures, which often struggle to model the complex spatial composition and semantic ambiguity in remote sensing images [7, 8]. These models typically adopt a single CNN encoder, such as ResNet [4], InceptionNet [8], or VGGNet [16], without exploring how encoder diversity impacts captioning performance [3]. However, single encoders are limited in capturing both global context and fine-grained object-level details due to constrained receptive fields and uniform feature hierarchies [4], which reduces the expressiveness of extracted visual features.
Decoder components exhibit similar bottlenecks. Many models use shallow decoders based on single-layer LSTM [3], GRU [8], or SVM [7], which often overlook hierarchical structure and intermediate semantics, leading to repetitive or oversimplified captions. Deeper decoders have been introduced in related contexts, but their exploration has not been as extensive for RSIC [1]. In addition, previous stacking architectures employ the final output layer alone [12], which ignores intermediate signals with complementary semantic information. The opportunity to use adaptive fusion schemes, which can dynamically aggregate multiple-layer decoder output, has not been thoroughly studied.
To address these drawbacks, we design a multi-stream RSIC architecture [26]. On the encoder side, two different CNNs are fused to exploit global semantics and local fine-scale spatial information. On the decoder side, we propose a stacked GRU sequence generator that makes use of the output of the intermediate layers with a locally weighted fusion mechanism. This enables the decoder to integrate multilevel linguistic cues and generate more informative and coherent descriptions. Experiments on several benchmark datasets demonstrate that the proposed approach is competitive and consistent with an improvement over single-stream baselines.
The major contributions of this work are as follows.
  • A fusion-based encoding strategy is proposed to improve image representation and feature diversity by integrating distinct features from two complementary CNN encoders.
@@ 1 Introduction
  • A novel multi-layer stacked GRU decoder with weighted averaging is proposed to improve semantic modeling of the input sequence.
The remainder of the paper is as follows. Section˜2 reviews the literature based on our pre-work. Section˜3 describes our proposed method. Section˜4 describes the experimental setup. Section˜5 presents the results. Finally, Section˜6 summarizes our main findings and provides directions for future work.
2 Related Work
Remote Sensing Image Captioning (RSIC) has seen significant advances in recent years, using deep learning techniques to generate textual descriptions of remote sensing images. Although natural image capture (NIC) has been extensively studied [31], RSIC remains a more specialized field with unique challenges due to its domain-specific characteristics [8]. The most widely adoptd framework in this domain is the encoder-decoder architecture [18], which forms the foundation of many state-of-the-art models.
Encoder-decoder architectures are commonly used in RSIC. Qu et al., [2016] [18] introduced a deep multimodal neural network for generating descriptions from remote sensing images. Lu et al., [2017] [16] performed a comparative analysis on existing models using datasets such as SYDNEY [18] and UCM [18], leading to the development of the RSICD dataset to address training data limitations. Das et al., [2024] [4] introduced an encoder-decoder framework with a text graph convolutional network and multilayer LSTM for automated remote sensing image captioning, significantly improving caption generation by enhancing semantic understanding and using a comparison-based beam search. Liu et al., [2025] [14] proposed a semantic-spatial feature fusion with dynamic graph refinement (SFDR) method for remote sensing image captioning, integrating multi-level feature representation and graph attention to improve the accuracy of generated descriptions. These methods significantly improved captioning accuracy, but still struggled with long-term dependencies and contextual consistency.
To overcome these limitations, attention mechanisms were introduced, significantly improving feature extraction and sequence modeling capabilities [29]. Sattar et al., [2024] [19] introduced the enhanced focal mechanism (EF) for image description generation, refining the focal point mechanism with multi-head attention and LSTM to improve correlation between image objects and improve subtitle quality. Wang et al., [2024] [24] presented a remote sensing image captioning model with sequential attention and flexible word correlation (SA-FWC), using VGG16 for feature fusion and LSTM for improved feature representation, achieving superior performance with minimal training sample pairs. Li et al. [2025] [12] proposed a feature refinement and rethinking of the attention framework for remote sensing image captioning, incorporating a refinement gate and rethinking LSTM to improve region focus and model discriminative contextual representations.
Despite these improvements, existing models still suffer from inefficiencies in encoder utilization and suboptimal feature aggregation in the decoder. Many studies focus on refining the decoder [7, 8], but optimal feature extraction remains an open challenge. To address these issues, our approach leverages a fusion-based encoding strategy that integrates multiple CNN architectures for enhanced feature representation. Additionally, we incorporate locally weighted stacking of decoders to retain hierarchical linguistic patterns effectively.
3 Proposed Method
@@
The overall architecture is illustrated in Figure˜1. We adopt an encoder-decoder framework, where the encoder ($F$) consists of two distinct CNNs whose outputs are concatenated and processed through a linear transformation layer ($L_{1}$). For decoding, a tokenized input sequence passes through an embedding layer ($E$) and then into a stack of Gated Recurrent Units (GRU) ($G$) to model temporal dependencies. The outputs from $L_{1}$ and $G$ are concatenated and passed through two additional linear layers ($L_{2}$ and $L_{3}$), with $L_{3}$ serving as a softmax layer to generate word probabilities.
The overall architecture of our work is shown in Figure˜1. We follow an encoder-decoder architecture, where the encoder ($F$) encompasses two separate CNNs, whose outputs are concatenated and processed by a linear transformation layer ($L_{1}$). In the decoder, an input sequence in tokens is stored in an embedding layer ($E$) followed by series of GRU cells ($G$) to capture temporal relationships. The outputs of $L_{1}$ and $G$ are concatenated together, which is then used as input to two more linear layers $L_{2}$, and another layer $L_{3}$ serving as a softmax to generate word probabilities.
3.1 Heterogeneous Encoding Strategy for Feature Diversity
Figure 2: Illustrative Example of Encoder Fusion in the Proposed Model
The proposed method integrates features extracted from two widely used CNN architectures. An illustrative example of the encoder fusion strategy is presented in Figure˜2. This study focuses on four prominent CNN architectures: ConvNext [15], ResNet [6], ResNext [28], and DenseNet [9]. These models have demonstrated strong feature extraction capabilities in RSIC [3, 2].
Table 1: Performance Comparison of Individual and Fusion-Based CNN Encoders on the SYDNEY Dataset
@@ 3.1 Heterogeneous Encoding Strategy for Feature Diversity
Table 2: Performance Comparison of Individual and Fusion-Based CNN Encoders on the UCM Dataset
@@
Table 3: Performance Comparison of Individual and Fusion-Based CNN Encoders on the RSICD Dataset
@@
The performance of various CNNs and their pairwise combinations is presented in Tables˜1, 2 and 3. As shown in the results, the combination of features from two distinct CNN architectures consistently improves performance compared to using a single CNN alone. This fusion does not add significant computational cost, as CNNs are used solely for feature extraction and remain frozen during training. Since the encoder parameters are not updated, the training remains efficient while benefiting from the diverse spatial and semantic information captured by multiple CNNs. Among the combinations, ConvNext and ResNet yield the best results due to their complementary properties. ConvNext effectively captures global spatial patterns, while ResNet retains fine-grained details through residual connections. Previous studies [2, 3] also highlight their superiority in RSIC tasks. Their fusion results in richer and more robust feature representations, making ConvNext and ResNet the chosen pair for all subsequent experiments in this study.
3.2 Deep Decoding Strategy for Semantic Modeling
In this paper, we have used the GRU [10] as a decoder. The GRU handles the flow of information through its gating mechanisms efficiently by retaining important temporal dependencies and being less computationally expensive. We have used a stacked GRU architecture in the decoder to improve the sequential modeling of the decoder. However, a trivial stacking (the output depends only on the last layer) may have the risk of discarding intermediate information and gradients [5]. To deal with this, we examine four methods of combining the models as follows.
3.2.1 Simple Stacking
The traditional stacking approach, called Simple Stacking (SS), directly utilizes the output of the final decoder layer as the final output of the system. The corresponding equation of SS is given below.
$O=D_{n}$ (1)
3.2.2 Concatenation Stacking
In Concatenation Stacking (CS), the outputs of all decoder layers are concatenated to form the final output, thus preserving a richer set of feature representations. The corresponding equation of CS is given below.
$O=D_{1}|D_{2}|...|D_{n}$ (2)
3.2.3 Global Weighted Stacking
In Global Weighted Stacking (GWS), a weighted average of the output of all decoder layers is calculated using a single global weight, ensuring a balanced contribution from each layer. The corresponding equation of GWS is given below.
$O=\sum_{i=1}^{n}\alpha_{i}{D_{i}}$ (3)
3.2.4 Local Weighted Stacking
In Local Weighted Stacking (LWS), a weighted average of the decoder output is calculated using element-specific weights, allowing for finer control over the contribution of each feature to the final output. The corresponding equation of LWS is given below.
$O=\sum_{i=1}^{n}W_{i}\odot D_{i}$ (4)
Here, $n$ denotes the number of decoders, and $D_{i}$ represents the output of the $i^{\text{th}}$ decoder, while $O$ is the final output of the stacked system. The symbol $|$ indicates the concatenation operation. The parameter $\alpha_{i}$ represents a scalar weight, while $W_{i}$ is a vector weight of the same size as $D_{i}$ for each $D_{i}$. Furthermore, $0\leq\alpha_{i}\leq 1$ with constraint $\sum_{i=1}^{n}\alpha_{i}=1$, and $0\leq W_{i}^{j}\leq 1$ with $\sum_{i=1}^{n}W_{i}^{j}=1$, where $W_{i}^{j}$ denotes the value of the characteristic $j^{\text{th}}$ of $W_{i}$. Both $\alpha=\{\alpha_{i}\mid i=1,\dots,n\}$ and $W=\{W_{i}\mid i=1,\dots,n\}$ are trainable parameters.
Table 4: Comparison of Different Decoder Stacking Methods on the SYDNEY Dataset
@@ 3.2.4 Local Weighted Stacking
Table 5: Comparison of Different Decoder Stacking Methods on the UCM Dataset
@@
Table 6: Comparison of Different Decoder Stacking Methods on the RSICD Dataset
@@
The impact of the depth of the decoder is evaluated in Tables˜4, 5 and 6, where the number of stacked GRUs (DC) is treated as a hyperparameter ranging from two to five. The number of decoders larger than five is excluded as they become computationally more expensive and yield lower returns in exchange.
Deeper stacking for SS and CS in the SYDNEY and UCM datasets(Tables˜4 and 5) has inconsistent trends that lead to performance degradation when the number of layers exceeds two. SS performs particularly worse than the single decoder baseline (NS), possibly due to vanishing gradients and the small size of the training data [16]. In contrast, the RSICD dataset (Table˜6) performs stably until the depth of three layers and declines beyond that, causing the performance insensitivity to the depth of the decoder in this dataset. From the results, we observe that among the stacking strategies, SS and CS are less effective. SS eliminates intermediate outputs, while CS introduces redundancy by treating all outputs equally. Weighted strategies offer a better balance: GWS uses global weights but lacks temporal flexibility, whereas LWS applies adaptive element-wise weights, enabling finer control over temporal dependencies. LWS consistently achieves the best results, with three stacked decoders yielding optimal performance across all datasets.
4 Experimental Setup
We use a conventional encoder-decoder architecture. Each GRU on the decoder stack ($G$) has an output size of $256$. Linear layers $L1$, $L2$, and $L3$ have output sizes of $256$, $512$, and one more than the vocabulary size, respectively. GELU is used as the activation function for $L1$ and $L2$, while $L3$ uses SoftMax. A dropout rate of $0.5$ is applied before $L1$, after the embedding layer ($E$), before the stack of linear layers ($SL$), and between consecutive GRUs to prevent overfitting. The embedding size is set to $256$.
@@ 4 Experimental Setup
The experiments were carried out using TensorFlow within a Docker environment with GPU support, on a host machine equipped with a NVIDIA RTX A6000 GPU with $50GB$ memory and $132GB$ RAM. During manuscript preparation, ChatGPT and Grammarly were used exclusively for rephrasing and language refinement, without generating original technical content.
4.1 Datasets Used
  • SYDNEY: SYDNEY [18] dataset consists of $613$ images, with 497 allocated for training, $58$ for testing, and $58$ for validation. It is derived from the Sydney dataset [32] and has been carefully curated to include seven distinct categories.
  • UCM: UCM [18] dataset consists of $2,100$ images, with $1,680$ allocated for training, $210$ for testing, and 21210 for validation. It is derived from the UC Merced Land Use dataset [30], which features 2121 land use image classes, each containing $100$ images.
  • RSICD: RSICD [16] dataset includes a substantial collection of $10921$ images, with 8,0348{,}034 designated for training, 1,0931{,}093 for testing and 1,0941{,}094 for validation. Sourced from a variety of platforms such as Google Earth [27], Baidu Map, MapABC, and Tianditu, this dataset encompasses $31$ distinct image classes.
In this study, we used the corrected versions of these datasets [4], which address common issues such as spelling mistakes, grammatical errors, and inconsistent dialect variations of English. We used the same train-validation-test split as defined in these datasets.
4.2 Performance metrics
Table 7: Experimental Results of Several RSIC Methods on the SYDNEY Dataset